### v1

# StyleReiser: Stylizing Video With Reinforced Structure Guide

Abstract
We introduce StyleReiser, an example-based video stylization method that transfers style from a given keyframe to the entire video sequence while maintaining visual consistency even in distant frames where the scene structure may change significantly. Unlike previous keyframe-based methods, our approach considers consistency with the prescribed style and maintains fidelity to new structural elements appearing in the target video sequence. This combination can significantly improve the quality of the stylized sequence without the need to add more correction keyframes. We also demonstrate that our approach can notably enhance the output of text-driven video stylization methods by suppressing their structural instability and enabling the user to perform custom edits on the generated keyframes. Moreover, due to its capability to perform inference in real-time, our technique can also be applied in interactive scenarios, such as consistently stylized video calls, which are difficult to achieve using text-driven approaches.
1 Introduction
Conditional video generation has received considerable attention recently. Given a combination of content and structure guidance, approaches that utilize large text-to-image diffusion models have achieved impressive results Yang et al. (2023); Ceylan et al. (2023); Geyer et al. (2024). However, the present video-to-video frameworks struggle with two significant issues: (i) poor temporal consistency of the stylized output, and (ii) insufficient fidelity to the input video (cf. Fig. 1d). These shortcomings can significantly impact the quality and usability of the generated content. Furthermore, to generate a single frame, considerable computational resources are required as a result of the use of diffusion models. Another important drawback of diffusion models is the limited control over stylized content. It is often difficult to come up with a text prompt that expresses exactly the intended artistic vision. In some cases manual intervention is inevitable, and it is not clear how to incorporate those additional edits. On the other hand, feed-forward video-to-video methods that are not dependent on diffusion models Jamriška et al. (2019); Texler et al. (2020b); Futschik et al. (2021) excel in semantically meaningful stylization, effectively preserving key features of the exemplar style. Users are granted full creative freedom with the ability to directly manipulate the style exemplar, bypassing the need for laborious refinement of text prompts, and have interactive feedback Texler et al. (2020b). However, when a new structural element appears in the input video, their results degrade considerably (cf. Fig. 1e), and additional stylized keyframes need to be provided. Those can be tedious to create manually or can introduce inconsistencies when generative methods are used for their preparation.
In this paper, we demonstrate that a key limitation of previous keyframe-based video stylization approaches lies in the fact that they are aiming to transfer the style but not explicitly trying to preserve structure of the target content. This omission may cause difficulties when significant changes occur in the frames farther away from the keyframe. To mitigate this drawback, we propose to relax on style guidance and focus more on the increased fidelity to novel structural elements appearing in the input video. This helps to significantly prolong the resistance to structural changes (cf. Fig. 1c) and thus eliminate the need to specify additional keyframes. Such an improvement brings benefits for both manually prepared keyframes where it saves time and effort of an artist as well as for synthetically generated ones where it prevents unwanted injection of structural inconsistencies caused by diffusion models. In addition, since our trained network can process the input video in real time, we can significantly reduce computational overhead in cases such as a video conference scenario (as was already demonstrated, e.g., in Futschik et al. Futschik et al. (2021)) where diffusion-based methods are difficult to apply.
(a) keyframe
(b) target frame
(c) our approach
(d) Ceylan et al. Ceylan et al. (2023)
(e) Futschik et al. Futschik et al. (2021)
Figure 1: An example of style transfer with reinforced structure guidance. We train a feed-forward neural network to transfer the style of a single stylized keyframe (a) to the remaining video frames (b) producing the output stylized frames (c). The keyframe (a) was created using text-driven video-to-video diffusion model of Ceylan et al. Ceylan et al. (2023). When the same model is applied on a different frame from the same video sequence (b), a significant inconsistency in structure can occur (d). The same holds for the result of state-of-the-art keyframe-based video stylization method Futschik et al. (2021). In contrast our approach (c) offers a stylized result that is faithful to the structural changes in the target video sequence and thus can deliver consistent output.
2 Related work
The origin of image and video stylization techniques can be traced back several decades. Early stylization approaches were typically based on hand-crafted algorithmic solutions that were restricted to a certain range of styles and specific target domains. For instance, Curtis et al. Curtis et al. (1997) runs a physically-based simulation to mimic the appearance of a watercolor media, Salisbury et al. Salisbury et al. (1997) produces painterly artworks automatically using a set of predefined brush strokes, while Praun et al. Praun et al. (2001) can generate brush strokes procedurally. Despite the impressive results these early stylization techniques produce, their main limitation lies in the fact that slight modification of an existing style or creation of a new one usually requires a significant effort and expertise.
[…]
Figure 2: An overview of our approach. Given images from a source domain $\mathbf{x}_{i}\in\mathcal{X}$, we optimize the network $f$ to produce images $\hat{\mathbf{y}}_{i}$ with a similar appearance as images from a target domain $\mathbf{y}_{k}$. The key loss $\mathcal{L}_{\text{key}}$ (1) encourages reconstruction of keyframes $\mathbf{y}_{k}$, the VGG loss $\mathcal{L}_{\text{VGG}}$ (2) ensures style consistency between frames and keyframes using Gram correlation matrices $g$ of extracted VGG network responses $v$ Gatys et al. (2016), and finally the cSDS loss $\mathcal{L}_{\text{cSDS}}$ (3) forces fidelity to structural elements present in the input video frames $\mathbf{x}_{i}$. cSDS loss requires a pre-trained ControlNet Zhang et al. (2023b) model consisting of diffusion model $d$ initialized by adding a random Gaussian noise $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ into the synthesized image $\hat{\mathbf{y}}_{i}$, a time step $t$, and a function $c$ transforming the input image $\mathbf{x}_{i}$ to a condition $\mathbf{c}$ (in this case Canny edge detector Canny (1986)).
To overcome this limitation, Hertzmann et al. Hertzmann et al. (2001) introduced the idea of image analogies. In their approach, the user can provide an example pair of unstyled and stylized images that specify the intended stylization analogy. The resulting image is then constructed by copying patches from a stylized exemplar so that the corresponding pixels in the unstyled patches match the patches in the target unstyled image. The framework of image analogies later proved to be a viable solution also for example-based video stylization Bénard et al. (2013); Jamriška et al. (2019) that can deliver temporally consistent sequences that faithfully preserve the user-specified visual style. A key limitation of those techniques is that they treat the target video as a guide for style transfer, and thus larger structural changes that may appear in the target domain are not taken into account. To overcome this limitation, the user needs to provide multiple consistently stylized keyframes, of which manual preparation can be labor intensive. When stylized keyframes are generated synthetically, it is, on the other hand, difficult to ensure their consistency.
Those limitations were addressed by Frigo et al. Frigo et al. (2016) and Gatys et al. Gatys et al. (2016) who perform stylization using only the style image and try to better respect the structural changes in the target domain. Frigo et al. search for the optimal mapping between the adaptively sized patches in the target image and patches in the style-exemplar while Gatys et al. iteratively optimizes the output image so that when fed into the VGG network Simonyan and Zisserman (2014) its responses correspond to VGG responses of the style exemplar and the target image. This approach inspired follow-up works Li et al. (2017); Kolkin et al. (2019) of which aim is to increase the faithfulness of the generated image to the style exemplar by employing more sophisticated loss functions. Chen et al. Chen et al. (2017) and Ruder et al. Ruder et al. (2018) later demonstrated how to extend the framework of Gatys et al. to example-based video stylization delivering temporally consistent sequences. Although those approaches are fully automatic and do not require preparation of a larger number of stylized keyframes, their artistic control over the final output is fairly limited. The transfer is usually not semantically meaningful and lacks faithfulness to the original artistic media.
To perform a semantically meaningful transfer while respecting structural changes in the target domain, image-to-image translation networks were proposed Johnson et al. (2016); Isola et al. (2017). However, these require a large amount of training data to work reliably. Only in some domain-specific scenarios, such as portrait stylization Futschik et al. (2019) training pairs can be generated automatically Fišer et al. (2017).
To mitigate the requirement for larger paired datasets, few-shot learning approaches Liu et al. (2019); Wang et al. (2019) and deformation-based approaches Siarohin et al. (2019b, a) were proposed. However, these methods require pretraining on large domain-specific datasets and thus are not applicable in the general case. Texler et al. Texler et al. (2020b) proposed a few-shot patch-based training strategy for which only a few stylized keyframes are necessary to deliver compelling video stylization results without the requirement of domain-specific pre-training. However, their method has limitations comparable to the image analogies approach of Jamriška et al. Jamriška et al. (2019), i.e., when new structural details appear in the target sequence, it is necessary to provide additional stylized keyframes. Although in the follow-up work of Futschik et al. Futschik et al. (2021) the amount of required keyframes decreased significantly, the newly appearing structural changes cause difficulties as the underlying approach is still predominantly focused on style preservation.
As an alternative approach to video stylization, unwrapping techniques have been proposed Rav-Acha et al. (2008); Kasten et al. (2021). In those approaches, input video frames are first projected onto a static atlas where edits can be performed at one snap and later transferred back to the original video domain. The quality of results is highly dependent on the quality of the generated unwrap. For small local edits, those techniques produce impressive results, but larger changes typically cause difficulties.
Recent approaches to video stylization utilize large pre-trained text-to-image diffusion models Rombach et al. (2022) and can edit videos globally using text prompts Khandelwal (2023); Yang et al. (2023); Zhang et al. (2023a); Chu et al. (2024); Geyer et al. (2024); Kim et al. (2024); Shin et al. (2024). However, the outputs of these techniques heavily depend on a particular version of the pre-trained text-to-image diffusion model, which may tend to produce unpredictable results that do not consistently reproduce structural changes in the target video sequence. This leads to a typical ubiquitous structural flicker that can be disturbing to the observer. Moreover, in addition to the text prompt, additional control over the stylization process is difficult to achieve, in contrast to keyframe-based methods where the user has full creative freedom Jamriška et al. (2019); Texler et al. (2020b); Futschik et al. (2021).
3 Method
A basic overview of our approach is given in Fig. 2. We train a consistent stylization operator (a modified variant of U-Net Ronneberger et al. (2015)) to stylize all frames of an input video sequence, given one (or possibly more) stylized keyframes. Those keyframes can be prepared manually, through a text-to-image generative model, or using a combination of both approaches, i.e., generating a stylized version of an input keyframe and then correcting it manually.
For the stylization operator $f$, we adopt the network architecture design proposed by Futschik et al. Futschik et al. (2019) that proved to be suitable for style transfer tasks Texler et al. (2020a); Futschik et al. (2021) allowing for reproduction of important high-frequency details, critical for generating complex and believable artistic styles.
$\mathbf{x}$
(a) Ceylan et al. Ceylan et al. (2023)
(b) Yang et al. Yang et al. (2023)
(c) Chu et al. Chu et al. (2024)
(d) Geyer et al. Geyer et al. (2024)
target frame
our approach
Figure 3: Results of our approach in comparison with the state-of-the-art in diffusion-based video stylization: The target video sequence (see a representative target frame $\mathbf{x}$) has been stylized using diffusion-based approaches (top row): (a) Ceylan et al. Ceylan et al. (2023), (b) Yang et al. Yang et al. (2023), (c) Chu et al. Chu et al. (2024), and (d) Geyer et al. Geyer et al. (2024). One frame from those stylized sequences was used as a keyframe (see small insets). The style of this keyframe has been propagated to the rest of the target sequence $\textbf{x}\in\mathcal{X}$ using our approach (bottom row). Note how our approach better preserves the structural details seen in the target frame. Also, see our supplementary video to compare consistency across the entire sequence. Diffusion-based approaches tend to suffer from notable structural flicker, whereas our approach keeps the structure consistent, yielding considerably more stable results.
Given two sets of images, we train $f$ to perform a consistent stylization on both sets. The first set $\mathcal{K}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\in\mathcal{X}\times\mathcal{Y}\mid i\in\mathbf{K}\}$ of keyframes contains $K$ tuples of original frames $\mathbf{x}\in\mathcal{X}$ and their stylized counterparts $\mathbf{y}\in\mathcal{Y}$. The symbol $\mathcal{X}$, $|\mathcal{X}|=N$, denotes a set of all input images, the symbol $\mathcal{Y}$, $|\mathcal{Y}|=K$, denotes a set of all stylized images, and the symbol $\mathbf{K}$, $|\mathbf{K}|=K$, denotes a set of frame indices to which a ground truth stylized version exists. The second set $\mathcal{Z}=\{\mathbf{x}_{i}\in\mathcal{X}\mid i\not\in\mathbf{K}\}$, $|\mathcal{Z}|=N-K$, consists of all frames of the input video to which there is no stylized counterpart. Both sets of images are used during the training step of our approach.
Let $f(\mathbf{x}_{i};\mathbf{\Phi})$ be the output of our neural network for the $i$-th input image and $\mathbf{\Phi}$ all the learnable parameters of the network. We use an $\ell_{2}$ loss
$\mathcal{L}_{\text{key}}(\mathcal{K};\mathbf{\Phi})=\frac{1}{|\mathbf{K}|}\sum_{i\in\mathbf{K}}\left\|f\left(\mathbf{x}_{i};\mathbf{\Phi}\right)-\mathbf{y}_{i}\right\|_{2}^{2}$ (1)
to encourage the $f$ operator towards reconstruction of the user-provided stylized keyframes.
In addition to the $\ell_{2}$ loss, we apply a distribution matching, VGG-based, style loss, computed as
$\begin{split}\mathcal{L}_{\text{VGG}}(\mathcal{K},\mathcal{Z};\mathbf{\Phi},\mathbf{\Psi})&=\\ =\frac{1}{N|L|}\sum_{i\not\in\mathbf{K}}\sum_{j\in\mathbf{K}}&\sum_{l\in\mathbf{L}}\left\|g(v(f(\mathbf{x}_{i};\mathbf{\Phi}),l;\mathbf{\Psi}))-g(v(\mathbf{y}_{j},l;\mathbf{\Psi}))\right\|_{2}^{2}\end{split}$ (2)
where $g(\mathbf{\cdot})$ computes the Gram correlation matrix Gatys et al. (2016), $v(\mathbf{\cdot},l;\mathbf{\Psi})$ extracts VGG network responses Simonyan and Zisserman (2014) at the layer $l$ with the VGG network parameters $\mathbf{\Psi}$, and $\mathbf{L}$ is a set of indices of all considered VGG layers. Distribution matching by minimizing differences between Gram matrices of neural VGG feature responses was originally proposed by Gatys et al. Gatys et al. (2016) for an optimization-based image stylization framework. Exactly as in their optimization framework, the parameters of the VGG network $\mathbf{\Psi}$ are frozen during the training of our operator $f$.
The last term of our operator loss is a variant of Score Distillation Sampling (SDS) proposed by Poole et al. Poole et al. (2022). SDS utilizes an off-the-shelf pre-trained text-to-image diffusion model to serve as a regularization technique in pursuit of realistically looking images. An image diffusion model is a generative model trained to invert the diffusion process, which progressively corrupts a natural image into images of statistically perfect Gaussian noise Ho et al. (2020). By removing small amounts of Gaussian noise with a fixed variance schedule $\bar{\alpha}_{t}$, an image is denoised. In their original work Poole et al. (2022), given a randomly sampled time step $t\sim\mathcal{U}(0,1)$, a random Gaussian noise $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, a noisy version of an image $\hat{\mathbf{y}}_{i}=f(\mathbf{x}_{i};\mathbf{\Phi})$, $\hat{\mathbf{y}}_{i,t}=\sqrt{\bar{\alpha}_{t}}\hat{\mathbf{y}}_{i}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon}$, and textual description $\mathbf{p}$, they show that a pre-trained text-to-image diffusion model $d(\hat{\mathbf{y}}_{i,t},\mathbf{p},t;\mathbf{\Theta)}$, parameterized by $\mathbf{\Theta}$, can be successfully utilized for optimization problems. In contrast to their work, we executed two key modifications of SDS. First, we made $t$ a constant. As shown in our experiments, the value of $t$ contributes to setting the level of fidelity to the input video structures. Second, we replaced the textual guidance by an image guidance. We set textual description $\mathbf{p}$ to an empty string and replaced a standard text-to-image diffusion model with a ControlNet Zhang et al. (2023b) additionally conditioned on $\mathbf{c}$ resulting in controlled Score Distillation Sampling (cSDS)
$\mathcal{L}_{\text{cSDS}}(\mathcal{Z};\mathbf{\Phi},\mathbf{\Theta})=\frac{1}{N}\sum_{i\not\in\mathbf{K}}\left\|d\left(\hat{\mathbf{y}}_{i,t},\mathbf{p},\mathbf{c},t;\Theta\right)-\epsilon\right\|_{2}^{2}.$ (3)
Note that the parameters $\mathbf{\Theta}$ of the “controlled” denoising diffusion model itself are fixed during the training of our operator. Also note that in practice, the back-propagation through the diffusion model is skipped to speed up the training. The condition $\mathbf{c}$ is the input image $\mathbf{x}$ transformed by a predefined function $c(\mathbf{x})$, e.g., Canny edge detector Canny (1986). The total loss for training the operator becomes
$\begin{split}\mathcal{L}(\mathcal{K},\mathcal{Z};\mathbf{\Phi},\mathbf{\Psi})=&\lambda_{k}\mathcal{L}_{\text{key}}(\mathcal{K};\mathbf{\Phi})+\\ +&\lambda_{v}\mathcal{L}_{\text{VGG}}(\mathcal{K},\mathcal{Z};\mathbf{\Phi},\mathbf{\Psi})+\\ +&\lambda_{c}\mathcal{L}_{\text{cSDS}}(\mathcal{Z};\mathbf{\Phi},\mathbf{\Theta}).\end{split}$ (4)
$\mathbf{x}$
(a) Ceylan et al. Ceylan et al. (2023)
(b) Yang et al. Yang et al. (2023)
(c) Chu et al. Chu et al. (2024)
(d) Geyer et al. Geyer et al. (2024)
target frame
our approach
Figure 4: Results of our approach in comparison with the state-of-the-art in diffusion-based video stylization (cont.): See Fig. 3 for a detailed explanation.
$\mathbf{x}$
(a) Ceylan et al. Ceylan et al. (2023)
(b) Yang et al. Yang et al. (2023)
(c) Chu et al. Chu et al. (2024)
(d) Geyer et al. Geyer et al. (2024)
target frame
our approach
Figure 5: Results of our approach in comparison with the state-of-the-art in diffusion-based video stylization (cont.): See Fig. 3 for a detailed explanation.
$\mathbf{x}$
(a) Ceylan et al. Ceylan et al. (2023)
(b) Yang et al. Yang et al. (2023)
(c) Chu et al. Chu et al. (2024)
(d) Geyer et al. Geyer et al. (2024)
target frame
our approach
Figure 6: Results of our approach in comparison with the state-of-the-art in diffusion-based video stylization (cont.): See Fig. 3 for a detailed explanation.
keyframe from Yang et al. Yang et al. (2023) $(\mathbf{x}_{r},\mathbf{y}_{r})$
custom edit $\mathbf{y}_{e}$
target frame $\mathbf{x}$
our method, stylization $\hat{\mathbf{y}}$
Figure 7: Custom edit of results generated by the diffusion-based method of Yang et al. Yang et al. (2023). Left to right: (a, b) keyframe $(\mathbf{x}_{r},\mathbf{y}_{r})$$\mathbf{y}_{r}$ is the result of Yang et al. Yang et al. (2023) conditioned with the text prompt “Galadriel, the royal Elf, silver-golden hair,” (c) custom edit $\mathbf{y}_{e}$ of the stylized keyframe $\mathbf{y}_{r}$, (d) target frame $\mathbf{x}$, (e) stylization $\hat{\mathbf{y}}$ produced by our method with a single keyframe $(\mathbf{x}_{r},\mathbf{y}_{e})$. A key advantage of our method is that it allows custom edits of videos stylized by diffusion-based methods, which provide only a limited textual control over the generated results.
target frame $\mathbf{x}$
4 keyframes
1 keyframe
(a) Ceylan et al. Ceylan et al. (2023)
(b) Jamriška et al. Jamriška et al. (2019)
(c) Texler et al. Texler et al. (2020b)
(d) Futschik et al. Futschik et al. (2021)
(e) our approach
Figure 8: Results of our approach in comparison with the state-of-the-art in keyframe-based video stylization: The diffusion-based method of Ceylan et al. Ceylan et al. (2023) has been used to generate a stylized sequence (a) from which $4$ keyframes were selected to perform video stylization using methods of Jamriška et al. Jamriška et al. (2019) (b) and Texler et al. Texler et al. (2020b) (c), and $1$ keyframe was selected for the method of Futschik et al. Futschik et al. (2021) (d) and our approach (e). Note how our approach better preserves the structural details seen in the target frame. In our supplementary video, it is also visible that our approach keeps the structure consistent, yielding notably more stable results.
target frame $\mathbf{x}$
4 keyframes
1 keyframe
(a) Geyer et al. Geyer et al. (2024)
(b) Jamriška et al. Jamriška et al. (2019)
(c) Texler et al. Texler et al. (2020b)
(d) Futschik et al. Futschik et al. (2021)
(e) our approach
Figure 9: Results of our approach in comparison with the state-of-the-art in keyframe-based video stylization (cont.): Diffusion-based method of Geyer et al. Geyer et al. (2024) has been used to generate the initial stylized sequence (a). See Fig. 8 for a detailed explanation.
target frame $\mathbf{x}$
4 keyframes
1 keyframe
(a) Yang et al. Yang et al. (2023)
(b) Jamriška et al. Jamriška et al. (2019)
(c) Texler et al. Texler et al. (2020b)
(d) Futschik et al. Futschik et al. (2021)
(e) our approach
Figure 10: Results of our approach in comparison with the state-of-the-art in keyframe-based video stylization (cont.): Diffusion-based method of Yang et al. Yang et al. (2023) has been used to generate the initial stylized sequence (a). See Fig. 8 for a detailed explanation.
4 Results and Comparison
We implemented our approach in PyTorch Paszke et al. (2019) using the AdamW Loshchilov and Hutter (2017) optimizer with fixed learning rate $3\cdot 10^{-5}$. Since the batch size was set to $1$, we used instance normalization Ulyanov et al. (2016) instead of batch normalization. We set $\lambda_{k}=1.0$ and $\lambda_{v}=100.0$. The parameters for $\mathcal{L}_{\text{cSDS}}$ (3) were selected experimentally and differ between the presented sequences, $\lambda_{c}\in\{10^{-5},10^{-6}\}$, $t\in\{20,28\}$. For score sampling purposes, we adopt the default noise scheduler of ControlNet v1.1 Zhang et al. (2023b) and the UniPC scheduler Zhao et al. (2024) with the number of steps set to $30$. The parameters of the VGG network and ControlNet were fixed and lineart conditioning was used (if not stated otherwise). The training was performed on a single NVIDIA A100 GPU with $40$ GB of RAM for $4$ hours and the model with the lowest total loss has been selected to produce the stylized sequences. The results are presented in Figures 1, 3, 4, 5, and 6. See also our supplementary material for additional results. We compare them with the output of recent diffusion-based methods Ceylan et al. (2023); Chu et al. (2024); Geyer et al. (2024); Yang et al. (2023) and keyframe-based approaches Futschik et al. (2021); Jamriška et al. (2019); Texler et al. (2020b).
4.1 Comparison with diffusion-based methods
Since diffusion-based methods only support textual guidance, for each sequence presented, we provide a textual description $\mathbf{p}_{\text{desc}}$ that describes the content of the sequence. We also provide the edit description $\mathbf{p}_{\text{edit}}$ that guides the stylization process. The resulting text prompt is a concatenation of both $\mathbf{p}=\mathbf{p}_{\text{edit}}\mathbf{p}_{\text{desc}}$, e.g., $\mathbf{p}_{\text{edit}}=`$hyperrealistic detailed oil painting of```` and $\mathbf{p}_{\text{desc}}=`$an old man with a white beard````. The description $\mathbf{p}_{\text{desc}}$ is used as an inversion prompt in the cases of Ceylan et al. Ceylan et al. (2023) and Geyer et al. Geyer et al. (2024). The target video sequences, represented in all figures by a single frame $\mathbf{x}$, were stylized using the mentioned diffusion-based methods (to run them, we used the configuration scripts provided by the authors in their implementations). A single frame from each stylized sequence was selected as a keyframe and its style was propagated throughout the target sequence $\mathbf{x}\in\mathcal{X}$ using our method. From the results presented in Figures 3, 4, 5, and 6 it is apparent that our approach maintains structural details better than diffusion-based approaches. See also our supplementary video that demonstrates structural consistency across the entire sequence contrasting the flicker common in diffusion-based methods.
A key advantage of our approach in contrast to diffusion-based techniques is that it enables the user to incorporate custom edits by manually modifying the output of the diffusion-based method and use it as a newly stylized keyframe for training. This option could be highly beneficial, especially in cases when it is difficult to find an accurate text prompt that precisely expresses the desired artistic vision. In Fig. 7, we show the results of a custom edit of an image $\mathbf{x}_{r}$ stylized by the method of Yang et al. Yang et al. (2023). The stylized image $\mathbf{y}_{r}$ was generated with the text prompt “Galadriel, the royal Elf, silver-golden hair.” Custom edits were performed on the image $\mathbf{y}_{r}$ producing the image $\mathbf{y}_{e}$. Our method was then trained with the keyframe $(\mathbf{x}_{r},\mathbf{y}_{e})$, producing the stylization $\hat{\mathbf{y}}$.
4.2 Comparison with keyframe-based methods
For comparison with keyframe-based approaches (see Figures 8, 9, and 10) we used sequences generated by diffusion-based methods and then $4$ keyframes of the stylized sequence were selected to train the methods of Jamriška et al. Jamriška et al. (2019) and Texler et al. Texler et al. (2020b), and $1$ keyframe was selected to train the method of Futschik et al. Futschik et al. (2021) and our method. In all results presented, our approach consistently maintains the structural details of the target frame $\mathbf{x}$ while faithfully replicating important style features of the stylized keyframe. See also our supplementary video to compare the structural consistency across the entire sequence.
5 Experiments
In this section, we present a set of experiments that provide better insight into how our approach performs in various settings. We first examine its convergence rate (Sec. 5.1), then we present an ablation study on our loss components (Sec. 5.2). We investigate the influence of parameters $\lambda_{c}$ and $t$ (Sec. 5.3) and conditioning prior $\mathbf{c}$ (Sec. 5.4). Finally, we compare the results of training with $\mathcal{L}_{\text{cSDS}}$ and without it using only the lineart guidance function $c(\mathbf{x})$ (Sec. 5.5).
5.1 Convergence rate
In the first experiment, we explore the optimization convergence speed of our method. The results presented in Fig. 11 indicate that a reasonably stylized output could be obtained after $6$ minutes of training, and training for more than $90$ minutes does not bring a significant improvement in stylization quality. Once the network has been trained, it is capable of performing a real-time stylization of a live video stream. This gives our method a key advantage over the diffusion-based video-to-video methods Ceylan et al. (2023); Chu et al. (2024); Geyer et al. (2024); Yang et al. (2023), which require at least $5$ seconds of computation per frame on a single NVIDIA A100 GPU.
1m
2m
3m
6m
10m
20m
45m
90m
Figure 11: Convergence speed of our method. Results captured at $1$, $2$, $3$, $6$, $10$, $20$, $45$, and $90$ minutes of training reveal the progressive improvement of the stylization. Reasonable results are retrieved after $6$ minutes of training, with no notable improvement beyond $90$ minutes on NVIDIA A100 GPU.
In Fig. 12, we compare stylization results of our method (Figures 12a–c, blue color plot) with Futschik et al. Futschik et al. (2021) (Figures 12d–f, orange color plot) after $30$, $60$, and $90$ minutes of training (plots are averages over $15$ different sequences; the border of the opaque area is the standard deviation). The approach of Futschik et al. Futschik et al. (2021) emphasizes the preservation of the original style. However, this emphasis results in a gradual loss of fine details within the structure of the target sequence, as illustrated by the zoom-in insets in Fig. 12. Note how the eye in Fig. 12f is directly copied from the stylized keyframe $\mathbf{y}_{k}$. A key advantage of our approach is that we explicitly retain these fine structural details, thus maintaining fidelity to the target frames $\mathbf{x}$.
00
$10^{-3}$
$10^{-2}$
$10^{-1}$
$30$
$60$
Minutes
Total    loss
$\mathbf{y}_{k}$
$\mathbf{x}_{k}$
$\mathbf{x}$
$90$
$120$
Figure 12: Comparative analysis of stylization results over training duration. In each of the first three columns, a stylized target frame $\mathbf{x}$ is shown after $30$, $60$, and $90$ minutes of training. Top row (a), (b), (c) and blue color: our method, bottom row (d), (e), (f) and orange color: Futschik et al. Futschik et al. (2021). The method of Futschik et al. Futschik et al. (2021) prioritizes the preservation of the original style. However, over time, this leads to a gradual loss of fine details in the structure of the target sequence (cf. zoom-in insets). A key advantage of our method is that we explicitly strive to retain these details, thereby maintaining their fidelity. The plots of the total loss are averaged over $15$ different sequences, the border of the opaque area depicts the measured standard deviation.
5.2 Ablation study on our loss components
In this experiment, we performed an ablation study to analyze the impact of individual loss terms within the optimization of our model. We systematically set multiplication constant of each of the three loss terms $\lambda_{k}$, $\lambda_{v}$, and $\lambda_{c}$ (4) to zero. The results are presented in Fig. 13 for the model with the lowest total loss after training on a GPU for $4$ hours. We specifically selected an input frame $\mathbf{x}$ with a significant appearance change with respect to the style exemplar $\mathbf{y}$ to highlight the ability of our method to produce results in the style of the exemplar even when new, previously unseen structures appear in the input frame. The $t=28$ and $\lambda_{\text{cSDS}}=5\cdot 10^{-6}$.
keyframe
target frame $\mathbf{x}$
$\mathcal{L}_{\text{VGG}}+\mathcal{L}_{\text{cSDS}}$
$\mathcal{L}_{\text{key}}+\mathcal{L}_{\text{cSDS}}$
$\mathcal{L}_{\text{key}}+\mathcal{L}_{\text{VGG}}$
$\mathcal{L}_{\text{key}}+\mathcal{L}_{\text{VGG}}+\mathcal{L}_{\text{cSDS}}$
Figure 13: An ablation study demonstrating the importance of individual terms in our objective function (4). A neural network is trained to transfer the style from the stylized keyframe $\mathbf{y}$ to the target frame $\mathbf{x}$. Each of the three terms $\mathcal{L}_{\text{key}}$ (1), $\mathcal{L}_{\text{VGG}}$ (2), and $\mathcal{L}_{\text{cSDS}}$ (3) is left out in the training. Leaving out $\mathcal{L}_{\text{key}}$ (a) causes the style transfer to be less semantically meaningful (see, e.g., white nose), excluding $\mathcal{L}_{\text{VGG}}$ (b) leads to complete style washout, and leaving out $\mathcal{L}_{\text{cSDS}}$ (c) results in poor replication of target frame structures. Only the combination of all three terms yield satisfactory results (d).
Excluding $\mathcal{L}_{\text{key}}$ from the loss in Fig. 13a results in an output that preserves structure of the target frame $\mathbf{x}$, however, the transfer is not semantically meaningful, i.e., colors from the stylized keyframe $\mathbf{y}$ are placed to improper locations in the target frame $\mathbf{x}$ (cf. white spots located on the nose and forehead).
When the term $\mathcal{L}_{\text{VGG}}$ is omitted in Fig. 13b, we get a result with an overall structure of the target frame $\mathbf{x}$ but with corrupted style details. This is caused by the fact that the term $\mathcal{L}_{\text{cSDS}}$ enforces the structure of the target frame $\mathbf{x}$ but not details of the stylized keyframe $\mathbf{y}$ that are maintained by the term $\mathcal{L}_{\text{VGG}}$.
When $\mathcal{L}_{\text{cSDS}}$ is omitted in Fig. 13c, the term $\mathcal{L}_{\text{VGG}}$ enforces the presence of the style details in the keyframe image $\mathbf{y}$, and the term $\mathcal{L}_{\text{key}}$ keeps the style transfer to be semantically meaningful. However, when previously unseen structures appear, such as new shapes of eyes, they are poorly reconstructed since the loss without term $\mathcal{L}_{\text{cSDS}}$ forces the network to reproduce only structures appearing in the stylized keyframe $\mathbf{y}$.
Only by combining the three loss terms together in Fig. 13d a result is produced that preserves the structural integrity of the target image $\mathbf{x}$ as well as incorporates the style features of the keyframe $\mathbf{y}$.
5.3 The influence of parameters $\lambda_{c}$ and $t$
In this experiment, we manipulate the parameter $t$ in $\mathcal{L}_{\text{cSDS}}$ (3) and the loss weight $\lambda_{c}$ (4). We trained the network for $4$ hours on a GPU and selected the network with the lowest total loss.
$t=16$
$t=20$
$t=24$
$t=28$
$\lambda_{c}=5\cdot 10^{-5}$
$\lambda_{c}=1\cdot 10^{-5}$
$\lambda_{c}=5\cdot 10^{-6}$
$\lambda_{c}=1\cdot 10^{-6}$
Figure 14: Reinforcement of input video structure as function of parameters $\lambda_{c}$ (4) and $t$ (3). The stronger $\lambda_{c}$, the more pronounced the structure of $\mathbf{x}$ is in the result. Note that the ControlNet used by the structural guidance term $\mathcal{L}_{\text{cSDS}}$ (3) uses a $30$ step denoising schedule. At $t=0$, the guidance input is pure Gaussian noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and at $t=29$, it contains minimal noise.
The results in Fig. 14 show that the stronger $\lambda_{c}$, the more pronounced the structure of $\mathbf{x}$ is in the result. Since the value of $\lambda_{c}$ modifies the strength of $\mathcal{L}_{\text{cSDS}}$ in the total loss term $\mathcal{L}$ (4), setting it to a higher value results in a stronger pronunciation of the input video structures in the output of our method. Another interesting dimension of structure reinforcement control strength is the parameter $t$. The ControlNet we utilize for the purpose of structure guidance uses a $30$ step denoising schedule. Every time the loss term $\mathcal{L}_{\text{cSDS}}$ is evaluated, a convex combination $\hat{\mathbf{y}}_{i,t}$ (3) of the output of our method and a Gaussian noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is performed. Then, the pre-trained ControlNet estimates the original noise $\epsilon$, and the residual between the estimated noise and $\epsilon$ is propagated through our method. This means that when $t=0$, ControlNet gets a pure Gaussian noise and when $t=29$, it gets the output of our network with only a little Gaussian noise present. According to this, we see that the lower the value of $t$, the stronger the focus is on the high-level structures compared to the higher values of $t$.
5.4 The influence of image conditioning prior $\mathbf{c}$
In our method, we use a ControlNet diffusion network Zhang et al. (2023b) to enforce the appearance of the structural elements of the target frame $\mathbf{x}$ in the stylized output of our network. We hypothesize that through sampling from the ControlNet network, we take advantage of both the structural conditioning provided by the guidance image $\mathbf{c}$ and the tendency of the network to guide the denoised samples toward the learned manifold of real images.
In this experiment, we investigate the suitability of four versions of pre-trained ControlNet models for stylization structural guidance. Each pre-trained ControlNet uses a different image guidance function $c$: (i) “lineart detection” neural model, (ii) “depth estimation” neural model, (iii) Canny edge detector Canny (1986), and (iv) “soft edge estimation“ neural model.
The results for $t=16$ and $t=28$ are presented in Fig. 15. Both depth and soft edge conditioning produce stylization that lacks details of the target frame $\mathbf{x}$. This appears to be caused by a very coarse structure of the two types of guidance, as seen in the guidance image $\mathbf{c}$. The Canny edge detector adds more signal to the ControlNet guidance, and the image guidance with the largest amount of signal is the lineart. Note that in applications where a fast training time is paramount, the quality of the lineart results might be traded for $\approx 10$ times faster Canny edge detector guidance. The depth guidance is in this sense impractical as it takes $\approx 10$ times more computational time than the lineart guidance, with the soft edge guidance somewhere in between the two.
In this experiment, we kept the $\lambda_{c}=5\cdot 10^{-6}$ and trained on a single NVIDIA A100 GPU for $4$ hours. We selected the model with the lowest total loss.
lineart
Canny
soft edge
depth
$\mathbf{c}$
$t=16$
$t=28$
target frame
stylized keyframe
Figure 15: Structure reinforcement as function of the image guidance type $\mathbf{c}$ and the number of diffusion time steps $t$. The best results are obtained using the lineart guidance that provides the ControlNet in $\mathcal{L}_{\text{cSDS}}$ (3) with the most detailed structural image and $t=28$.
5.5 Direct optimization with the image conditioning prior
$\mathcal{L}_{\text{cSDS}}$
$\mathcal{L}_{\text{lineart}}$
$\xrightarrow{\hskip 50.6459pt}$ increasing strength of loss term $\xrightarrow{\hskip 50.6459pt}$
Figure 16: Comparison of a stylization results obtained by the optimization with the lineart loss term $\mathcal{L}_{\text{lineart}}$ (5) and with the lineart-guided $\mathcal{L}_{\text{cSDS}}$ term (3). Increasing the strength of the loss term $\mathcal{L}_{\text{lineart}}$ does not enforce the structural elements from the target frame to the stylized output as in the case of the $\mathcal{L}_{\text{cSDS}}$ term.
In our experiments, we show that our proposed loss term $\mathcal{L}_{\text{cSDS}}$ (3) enforces structures from the target frame $\mathbf{x}$ to the stylized output of our method. The purpose of this experiment is to show that the $\mathcal{L}_{\text{cSDS}}$ loss built on the ControlNet model cannot be replaced by a simpler loss that uses the lineart image conditioning model $c(\mathbf{x})$, which is already part of ControlNet, where it provides the guidance image $\mathbf{c}$. In Fig. 16 we show that replacing the lineart-guided $\mathcal{L}_{\text{cSDS}}$ with the lineart term
$\mathcal{L}_{\text{lineart}}(\mathcal{Z};\mathbf{\Phi},\mathbf{\Theta}_{c})=\frac{1}{N}\sum_{i\not\in\mathbf{K}}\left\|c\left(\hat{\mathbf{y}}_{i};\mathbf{\Theta}_{c}\right)-c\left(\mathbf{x}_{i};\mathbf{\Theta}_{c}\right)\right\|_{2}^{2},$ (5)
where $\mathbf{\Theta}_{c}$ are the parameters of the model $c(\mathbf{x};\mathbf{\Theta}_{c})$ producing the lineart control image $\mathbf{c}$, does not enforce the structures from the target frame $\mathbf{x}_{i}$ to the stylized output $\hat{\mathbf{y}}_{i}$. In this experiment, we trained on a single NVIDIA A100 GPU for $4$ hours, and we selected the model with the lowest total loss.
6 Limitations
Although our approach represents an improvement over the current state-of-the-art in diffusion-based as well as keyframe-based video stylization methods, we have identified certain limitations in its application.
In style transfer methods, there exists a delicate balance between maintaining fidelity to the features of the style exemplar images and preserving the structural characteristics present in the content that is being stylized. Current state-of-the-art in keyframe-based video stylization mainly emphasizes fidelity to the original features in the style image. In our approach we aim to produce the stylized content aligned both with the style exemplar as well as with the structural characteristics of the target video sequence. Although we enable the user to find a right balance between style and structure using parameters $\lambda_{c}$ and $t$, we acknowledge that in some situations decreasing the fidelity to the style features may be perceived as a potential limitation.
Our method achieves comparable training time as the approach of Futshik et al. Futschik et al. (2021) which is relatively long and thus can limit the possibility to perform interactive updates that are available, e.g., in the method of Texler et al. Texler et al. (2020b). To address this drawback, in future work we plan to combine Texler et al.’s rapid patch-based training strategy with the calculation of VGG loss in a full-frame setting.
In our proposed workflow, a user is expected to select a keyframe that will be used for training. Although certain guidelines can be followed, a mechanism that would enable automatic selection of suitable keyframes could simplify the preparation phase.
7 Conclusion
We presented a novel keyframe-based video stylization method that maintains visual coherence across distant frames by adapting to changes in scene structure and adhering to important characteristics of a prescribed visual style. This capability not only enhances the quality of stylized sequences, it also eliminates the need for additional correction keyframes, which can be tedious to prepare or inconsistent when created using diffusion-based methods. Our approach serves as a viable alternative to text-based techniques because it mitigates their structural instability and allows users to make additional modifications to the generated keyframes. Moreover, its real-time inference capability enables interactive applications such as consistently stylized video calls, which are difficult to achieve using diffusion-based approaches.